OpenStudy (anonymous):

Find all absolute and relative extreme values of the function. f(x) = 2x/x - 4 on (-2,2) Choose the correct answer.

OpenStudy (anonymous):

A. The absolute maximum is f(4) = ∞ and the absolute minimum is f(2) = -2. B. The absolute maximum is f(2) = -2 and the absolute minimum is f(-2) = 2/3 . C. The absolute maximum is f(-2) = 2/3 and the absolute minimum is f(2) = -2. D. The absolute maximum is f(-2) = 2/3 and the absolute minimum is f(4) = -∞ .

OpenStudy (anonymous):

@electrokid can u help?

OpenStudy (anonymous):

are you there @electrokid

OpenStudy (anonymous):

1) find f'(x) 2) make f'(x) =0 and solve for "x" 3) find f''(x) 4) use second drivative test

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the derivatives only give you the relative extrema. find f(a) and f(b) for absolute extrema

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got A is that correct?

OpenStudy (anonymous):

are u there?

OpenStudy (anonymous):

but "4" is not in your domain!! \(x\in(-2,2)\)

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

ok i got C

OpenStudy (anonymous):

is it correct?

OpenStudy (anonymous):

good job